OpenStudy (anonymous):

This is the conics when I am given with the equation like 16y^2-x^2+2x+64y+63=0, I know how to do everything else, but when I complete the square and make it into standard form, the right side becomes 0, not 1. In this case, what should I do?

OpenStudy (anonymous):

I'll take a stab at it

OpenStudy (anonymous):

\[16y ^{2}-x ^{2}+2x+64y+63=0\]

OpenStudy (anonymous):

\[16\left( y ^{2} +4y+4\right)-\left( x ^{2}+2x+1 \right)=-63+64+1\]

OpenStudy (anonymous):

yes I got that :(

OpenStudy (anonymous):

\[\frac{ 16 }{ 2 }(y+2)^{2}-\frac{ (x+1)^{2} }{ 2 } = 1\]

OpenStudy (anonymous):

ohoh hold on

OpenStudy (anonymous):

how did you get +1 ?

OpenStudy (anonymous):

\[\frac{ (y+2)^{2} }{ \frac{ 2 }{ 16 }}-\frac{ (x+1)^{2} }{ 2 } = 1\]

OpenStudy (anonymous):

sorry for −63+64+1, how did you get 1? isn't it -1?

OpenStudy (anonymous):

-63+64 = 1

OpenStudy (anonymous):

then 1+1 = 2

OpenStudy (anonymous):

then divide both sides by 2

OpenStudy (anonymous):

okay Thank you! :)

OpenStudy (anonymous):

You're welcome